OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (dumbcow):

V = 1/3*pir^2*h 300 = 1/3*pi*49*h 900 = pi*49h (900/49pi) = h